### v1

@@ 2 Paley ETFs
We will apply the tools from Sections 1.2 and 1.3 to combinatorially and geometrically characterize one infinite class of ETFs.
We begin by reviewing some basic facts about finite fields. If $q=p^{s}$ for some prime $p$ and $\mathbb{F}_{q}$ is the finite field of size $q$, then $\mathbb{F}_{q}^{\times}=\mathbb{F}_{q}\backslash\{0\}$ is a cyclic group under multiplication. We denote the quadratic residues and non-residues as
$\operatorname{QR}(\mathbb{F}_{q})$ =$=\left\{\,a^{2}\,\,\middle|\,\,a\in\mathbb{F}_{q}^{\times}\,\right\}$
$\operatorname{NR}(\mathbb{F}_{q})$ $=\mathbb{F}_{q}^{\times}\backslash\operatorname{QR}.$
Thus, $\mathbb{F}_{q}=\operatorname{QR}(\mathbb{F}_{q})\cup\operatorname{NR}(\mathbb{F}_{q})\cup\{0\}$. When the underlying field is clear from context, we will often simply write $\operatorname{QR}$ and $\operatorname{NR}$. Note that $\operatorname{QR}$ is thus a subgroup of the multiplicative cyclic group $\mathbb{F}_{q}^{\times}$ of index $2$, meaning it has size $(q-1)/2$ and is also cyclic. When $q\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$; that is, the non-quadratic residues are precisely the negations of the quadratic residues. This fact implies the following (see, e.g. [CD07]).
Theorem 2.1.
When $q\equiv 3\bmod 4$, $\operatorname{QR}(\mathbb{F}_{q})$ forms a difference set in the additive group of $\mathbb{F}_{q}$.
@@
i.e., precisely the difference set in $\mathbb{Z}_{7}$ (i.e., the additive group of $\mathbb{F}_{7}$) in Example 1.9.
Thus, it follows from Theorem 1.10 that if $q\equiv 3\bmod 4$, $F$ is the Fourier transform of the additive group of $\mathbb{F}_{q}$, and $\Phi$ is the $(q-1)/2\times q$ submatrix of $F$ consisting of the rows indexed by $\operatorname{QR}(\mathbb{F}_{q})$, then $\Phi$ must be an ETF, which we call a Paley ETF. If $q\equiv 3\bmod 4$, we will denote by $\Phi_{q}$ this corresponding ETF.
@@
The smallest non-trival Paley ETF is $\Phi_{7}$. Namely, let $\zeta$ be a primitive $7$th root of unity. Then
@@
is one ordering of the $(3,7)$-ETF which is the Paley ETF for $\mathbb{F}_{7}$. Note that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ and thus $\operatorname{NR}(\mathbb{F}_{7})=\{3,5,6\}$. Further, other than the 00th column, half of the columns have exponents that are $\operatorname{QR}$ and half have $\operatorname{NR}$. Finally, note that (cf. Example 1.12)
@@
The properties seen in Example 2.3 generalize. First, we define two special operators.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $a$ be a generator of $\operatorname{QR}(\mathbb{F}_{p})$. The Paley modulation $M_{\{p\}}$ is the $(p-1)/2\times(p-1)/2$ diagonal matrix
$M_{\{p\}}=\begin{pmatrix}\zeta^{a^{0}}&0&\ldots&0\\ 0&\zeta^{a^{1}}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\zeta^{a^{(p-3)/2}}\end{pmatrix}.$
@@
$T_{d}=\begin{pmatrix}0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ 0&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&0\end{pmatrix}$
For a given $p$, the sizes of interest will be $M_{\{p\}}$ and $T_{(p-1)/2}$; thus, when clear from the context, we will often simply write $M$ and $T$.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $a$ be a generator of $\operatorname{QR}(\mathbb{F}_{p})$. Then the Paley ETF $\Phi_{p}$ may be constructed as
@@
This trivially follows from Example 1.12 and the fact that
$\operatorname{QR}(\mathbb{F}_{p})=\left\{\,a^{\ell}\,\,\middle|\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}.$
Note that, unlike in Example 2.3, the ordering of rows indexed using the construction above are not necessarily the ordering as a subsequence of $(1,\,\ldots,\,p-1)$. Proposition 2.5 also implies that other than $\mathbbm{1}$, $(p-1)/2$ vectors in $\Phi_{p}$ have entries equal to $\zeta$ raised to some permutation of the set of the $\operatorname{QR}$ and $(p-1)/2$ vectors entries which have $\operatorname{NR}$ as the exponents.
In order to use Theorem 1.14, we need to characterize the triple products of the Paley ETF vectors.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Further let $\zeta$ be the primitive $p$th root of unity and $a$ the generator of $\operatorname{QR}(\mathbb{F}_{p})$ used to generate $\Phi_{p}$. Finally, set
$\alpha=\sum_{\ell=0}^{(p-3)/2}\zeta^{a^{\ell}}.$
@@
$\operatorname*{TP}(j,k,\ell)=\left\{\begin{array}[]{lr}\alpha^{3};&j-k,k-\ell,\ell-j\in\operatorname{QR}\\[2.0pt] \left|\alpha\right|^{2}\alpha;&\textrm{ two of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \left|\alpha\right|^{2}\overline{\alpha};&\textrm{ one of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \overline{\alpha}^{3};&\ j-k,k-\ell,\ell-j\in\operatorname{NR}\end{array}\right.$
@@
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\sum_{\ell=0}^{(p-3)/2}\zeta^{ja^{\ell}}\zeta^{-ka^{\ell}}=\sum_{\ell=0}^{(p-3)/2}\zeta^{(j-k)a^{\ell}}=\left\{\begin{array}[]{lr}(p-1)/2&j=k\\ \alpha&j-k\in\operatorname{QR}\\ \overline{\alpha}&j-k\in\operatorname{NR}\end{array}\right.$
since $(j-k)a^{\ell}\in\operatorname{QR}$ if and only if $j-k\in\operatorname{QR}$ and also $\operatorname{NR}=-\operatorname{QR}$. So, for distinct $j,k,\ell$, the desired result immediately follows. ∎
We prove one more lemma before getting to the main result.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Let $\alpha$ be as in Lemma 2.6. Define $\Phi_{\operatorname{QR}}$ to be any $(p-1)/2\times(p-1)/2$ matrix consisting of the columns of $\Phi_{p}$ corresponding to the quadratic residues. Similarly define $\Phi_{\operatorname{NR}}$ for the non-quadratic residues. Then the row sums of $\Phi_{\operatorname{QR}}$ are $\alpha$, $\Phi_{\operatorname{QR}}$ is invertible, and the row sums of $\Phi_{\operatorname{QR}}^{-1}$ are $1/\alpha$. Similarly, the row sums of $\Phi_{\operatorname{NR}}$ are $1/\overline{\alpha}$, $\Phi_{\operatorname{NR}}$ is invertible, and the row sums of $\Phi_{\operatorname{NR}}^{-1}$ are $1/\overline{\alpha}$.

### v2

@@ 2 Paley ETFs
We will apply the tools from Sections 1.2 and 1.3 to combinatorially and geometrically characterize one infinite class of ETFs, the Paley (difference set) ETFs. These ETFs are generated using the structure of finite fields. Thus, we begin in Section 2.1 by reviewing some basic facts about finite fields. For each prime power $q\equiv 3\bmod 4$, there is a Paley ETF $\Phi_{q}$. We will prove in Sections 2.22.4 that $\operatorname*{Sym}_{v}(\Phi_{q})=\operatorname*{Sym}_{\ell}(\Phi_{q})$ act $2$-homogeneously but not $2$-transitively on the vectors and the lines, respectively. The Paley ETFs thus have a bender which is a BIBD. In Section 2.2, we prove the result (Theorem 2.10) when $q$ is prime as the proof requires less machinery and gives a flavor of the arguments for the more general case. This result also explicitly gives the unitary matrices which are vector automorphisms. Then in Section 2.3, we provide an alternate proof of Theorem 2.10 that filters through Theorem 1.14. This proof does not yield the explicit vector automorphisms but yields techniques which may be useful for other classes of ETFs. A quick example of the technique being applied to another infinite class of ETFs (Gabor-Steiner) is also given. Finally, in Section 2.4, we prove the result (Theorem 2.19) for all prime powers $q\equiv 3\bmod 4$.
2.1 Finite Fields
[…]
$\operatorname{QR}(\mathbb{F}_{q})$ $=\left\{\,x^{2}\,\,\middle|\,\,x\in\mathbb{F}_{q}^{\times}\,\right\}$
$\operatorname{NR}(\mathbb{F}_{q})$ $\QR.\displaystyle=\mathbb{F}_{q}^{\times}\backslash\operatorname${QR}.
Thus, $\mathbb{F}_{q}=\operatorname{QR}(\mathbb{F}_{q})\cup\operatorname{NR}(\mathbb{F}_{q})\cup\{0\}$. When the underlying field is clear from context, we will often simply write $\operatorname{QR}$ and $\operatorname{NR}$. Note that $\operatorname{QR}$ is thus a subgroup of the multiplicative cyclic group $\mathbb{F}_{q}^{\times}$ of index $2$, meaning it has size (q1)/2(q-1)/2 and is also cyclic and generated by a square of a primitive element.
@@ 2.1 Finite Fields
i.e., precisely the difference set in $\mathbb{Z}_{7}$ (i.e., the additive group of $\mathbb{F}_{7}$) in Example 1.9. Note further that $3$ and $5$ are primitive elements, i.e., $3>=<5>=\mathbb{F}_{7}^{\times}$ and that $2=3^{2}$ and $4=5^{2}$ generate $\operatorname{QR}(\mathbb{F}_{7})$.
For fields of prime order $p$, the structure is relatively simple. The additive group $(\mathbb{F}_{p},+)$ and the multiplicative group $(\mathbb{F}_{p}^{\times},\cdot)$ are both cyclic. For $q=p^{s}$ with $s>1$, the additive group $(\mathbb{F}_{q},+)$ is no longer cyclic. Further, we can think of $\mathbb{F}_{q}$ as both a field extension of $\mathbb{F}_{p}$ and a vector space over $\mathbb{F}_{p}$. We will need to leverage both types of structure in our analysis.
First, we represent $\mathbb{F}_{q}$ as a field extension. We pick a monic irreducible (and thus separable since $\mathbb{F}_{p}$ is a finite field) polynomial $\pi\in\mathbb{F}_{p}[X]$ of degree $s$. Then $\mathbb{F}_{q}$ is the Galois extension $\mathbb{F}_{q}\cong\mathbb{F}_{p}(\alpha)\cong\mathbb{F}_{p}[X]/(\pi(X))$, where $\pi(\alpha)=0$. Since $\mathbb{F}_{q}$ is also congruent to the splitting field of $X^{q}-X$, we can (and will) choose $\pi$ so that $\alpha$ is a primitive element and $\alpha^{2}$ generates $\operatorname{QR}$.
We may view the vector space structure of $\mathbb{F}_{q}$ as coordinate vectors in $\mathbb{F}_{p}^{s}$ via the primitive element basis $\alpha^{0},\,\alpha^{1},\,\ldots\alpha^{s-1}$. The Frobenius automorphism $x\mapsto x^{p}$ generates the Galois group of automorphisms. A basis which is the orbit of an element under the Galois group, i.e., $\left\{\,x^{p^{i}}\,\,\middle|\,\,i\in\{0,1,\ldots,s-1\}\,\right\}$ for particular $x\in\mathbb{F}_{q}$ is called a normal basis.
Example 2.2.
Consider $\pi(x)=x^{3}+2x+1\in\mathbb{F}_{3}[X]$. Then $\pi$ is irreducible over $\mathbb{F}_{3}$ since $0^{3}+2(0)+1,1^{3}+2(1)+1,2^{3}+2(2)+1\neq 0$. Let $\alpha$ denote a root of $\pi$. Then
$\begin{array}[]{ccccc}\alpha^{0}=1,&\alpha^{1}=\alpha,&\alpha^{2}=\alpha^{2},&\alpha^{3}=2+\alpha,&\alpha^{4}=2\alpha+\alpha^{2},\\ \alpha^{5}=2+\alpha+2\alpha^{2},&\alpha^{6}=1+\alpha+\alpha^{2},&\alpha^{7}=2+2\alpha+\alpha^{2},&\alpha^{8}=2+2\alpha^{2},&\alpha^{9}=1+\alpha,\\ \alpha^{10}=\alpha+\alpha^{2},&\alpha^{11}=2+\alpha+\alpha^{2},&\alpha^{12}=2+\alpha^{2},&\alpha^{13}=2,&\alpha^{14}=2\alpha,\,\ldots\end{array}$
where $\alpha^{13+k}=2\alpha^{k}$ for $k\in\{0,\ldots,13\}$. That is, $\alpha$ is a primitive element, $\{\alpha^{0},\alpha^{1},\alpha^{2}\}$ is a primitive element basis for $\mathbb{F}_{27}\cong\mathbb{F}_{3}(\alpha)$, and $\operatorname{QR}(\mathbb{F}_{27})=<\alpha^{2}>$. We also note that the orbit of $\alpha^{2}$ under the Galois group forms a normal basis. Namely,
$\alpha^{2},\quad\left(\alpha^{2}\right)^{3}=\alpha^{6}=1+\alpha+\alpha^{2},\quad\left(\alpha^{2}\right)^{9}=\alpha^{18}=2\alpha^{5}=1+2\alpha+\alpha^{2}$
form a basis for $\mathbb{F}_{27}$. One way to see this is by inspecting the coordinate vectors with respect to the primitive element basis:
$\begin{pmatrix}0&0&1\end{pmatrix}^{\top},\quad\begin{pmatrix}1&1&1\end{pmatrix}^{\top},\quad\begin{pmatrix}1&2&1\end{pmatrix}^{\top}$.
We denote the vector space of $s\times s$ matrices with entries in a field $\mathbb{F}$ as $\mathbb{F}^{s\times s}$. Let $C\in\mathbb{F}_{p}^{s\times s}$ be the companion matrix of $\pi$. Then the group generated by $C$ is isomorphic to $\mathbb{F}_{q}^{\times}$ via the map $C\mapsto\alpha$.
Example 2.3.
Consider the construction of $\mathbb{F}_{27}$ from Example 2.2. The companion matrix of $\pi$ in $\mathbb{F}_{3}^{3\times 3}$ is
$C=\begin{pmatrix}0&0&2\\ 1&0&1\\ 0&1&0\end{pmatrix}$.
Let $e_{1}$ be the first standard basis vector (i.e., $e_{1}$ represents the multiplicative identity in $\mathbb{F}_{27}$). Then, $Ce_{1}$ represents $\alpha$ using coordinates from the primitive basis $\{1,\alpha,\alpha^{2}\}$ and further
$\mathbb{F}_{27}^{\times}=\left\{\,C^{k}e_{1}\,\,\middle|\,\,k\in\{0,\ldots,26\}\,\right\}\quad\textrm{and}\quad\operatorname{QR}(\mathbb{F}_{27})=\left\{\,C^{2k}e_{1}\,\,\middle|\,\,k\in\{0,\ldots,12\}\,\right\}.$
Also, $\left\{\,C^{2k}e_{1}\,\,\middle|\,\,k\in\{1,3,9\}\,\right\}$ forms a normal basis for $\mathbb{F}_{27}$.
When $q\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$; that is, the quadratic non-residues are precisely the negations of the quadratic residues. This fact implies the following (see, e.g. [CD07]).
Theorem 2.4.
When $q\equiv 3\bmod 4$, $\operatorname{QR}(\mathbb{F}_{q})$ forms a (q,(q1)/2,(q3)/4)(q,(q-1)/2,(q-3)/4)-difference set in the additive group of $\mathbb{F}_{q}$.
Thus, it follows from Theorem 1.10 that if $q\equiv 3\bmod 4$, $F$ is the Fourier transform of the additive group of $\mathbb{F}_{q}$, and $\Phi$ is the $(q-1)/2\times q$ submatrix of $F$ consisting of the rows indexed by $\operatorname{QR}(\mathbb{F}_{q})$, then $\Phi$ must be an ETF, which we call a Paley ETF. If $q\equiv 3\bmod 4$, we will denote by $\Phi_{q}$ this corresponding ETF. We will fix an ordering of the rows of $\Phi_{q}$ as $\alpha^{0},\alpha^{2},\ldots,\alpha^{q-3}$ for a primitive element $\alpha$. This will allow the explicit definition of matrices in $\operatorname{U}((q-1)/2)$ which are the vector automorphisms. The ordering of the columns is less important than the parameterization, as we just need to characterize the induced permutations of $\mathbb{F}_{q}$.
@@
The smallest nontrivial Paley ETF is $\Phi_{7}$. Namely, let $\zeta$ be a primitive $7$th root of unity. Then
@@
is one ordering of the $(3,7)$-ETF which is the Paley ETF for $\mathbb{F}_{7}$. Note that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ and thus $\operatorname{NR}(\mathbb{F}_{7})=\{3,5,6\}$. Further, other than the 0th column, half of the columns have exponents that are $\operatorname{QR}$ and half have $\operatorname{NR}$. Finally, note that (cf. Example 1.12)
@@
The ordering of the rows in this example corresponds to the choice of primitive element $3$, i.e.,
$\left(3^{2(0)},\,3^{2(1)},\,3^{2(2)}\right)=\left(2^{0},\,2^{1},\,2^{2}\right).$
Consider the following bilinear form on $\mathbb{F}_{p}^{s}$:
$\left\langle\begin{pmatrix}a_{0}\\ a_{1}\\ \vdots\\ a_{s-1}\end{pmatrix},\begin{pmatrix}b_{0}\\ b_{1}\\ \vdots\\ b_{s-1}\end{pmatrix}\right\rangle_{p}=\sum_{i=0}^{s-1}a_{i}b_{i}.$
Let $p$ be prime and $s\geq 0$. Further let $\zeta$ be a primitive $p$th root of unity. Up to possible reorderings of the rows and columns, the Fourier matrix of $(\mathbb{F}_{p}^{s},+)$ has entries $(\zeta^{\left\langle x,y\right\rangle_{p}})_{x,y\in\mathbb{F}_{p}^{s}}$. This follows immediately from the Kronecker product construction in Definition 1.7.
Example 2.6.
Continuing Examples 2.2 and 2.3, let $\zeta$ be a primitive third root of unity so that the Paley ETF $\Phi_{27}$ has entries
$\left\{\,\zeta^{\left\langle x,y\right\rangle_{p}}\,\,\middle|\,\,x\in\operatorname{QR},\,y\in\mathbb{F}_{27}\,\right\}.$
Since $Ce_{1}$ is a primitive element and $C^{2}e_{1}$ generates the quadratic residues, we may parameterize each column except the one indexed by $y=0$ (which has all entries equal to 11) as
$\left\{\,\zeta^{\left\langle x,y\right\rangle_{p}}\,\,\middle|\,\,x\in\operatorname{QR},\,y\in\mathbb{F}_{27}^{\times}\,\right\}=\left\{\,\zeta^{\left\langle C^{2k}e_{1},C^{\ell}e_{1}\right\rangle_{p}}\,\,\middle|\,\,k\in\{0,\ldots,12\},\,\ell\in\{0,\ldots,25\}\,\right\}.$
2.2 Paley ETFs over Fields of Prime Order
In this section, we will explicitly give the vector automorphisms of $\Phi_{p}$ for $p\equiv 3\bmod 4$ prime and show that the vector and line symmetry groups act $2$-homogeneously but not $2$-transitively. The line automorphism group is generated by the following operators.
@@ 2.2 Paley ETFs over Fields of Prime Order
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $\alpha$ be a primitive element of $\mathbb{F}_{p}$. The (cyclic) Paley modulation $M_{\{p\}}$ is the $(p-1)/2\times(p-1)/2$ diagonal matrix
$M_{\{p\}}=\begin{pmatrix}\zeta^{\alpha^{0}}&0&\ldots&0\\ 0&\zeta^{\alpha^{2}}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\zeta^{\alpha^{p-3}}\end{pmatrix}$.
@@
$T_{d}=\begin{pmatrix}0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ 0&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&0\end{pmatrix}$.
For a given $p$, the sizes of interest will be $M_{\{p\}}$ and $T_{(p-1)/2}$; thus, when clear from the context, we will often simply write $M$ and $T$. As in Example 2.5, we may characterize Paley ETFs over prime-order fields as the orbit under the cyclic group of Paley modulations.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Then the Paley ETF $\Phi_{p}$ may be constructed as
@@
Let $\alpha$ be the primitive element of $\mathbb{F}_{p}$ used to construct $M_{\{p\}}$. The result trivially follows from Example 1.12 and the fact that
$\operatorname{QR}(\mathbb{F}_{p})=\left\{\,\alpha^{2\ell}\,\,\middle|\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}.$
Note that, unlike in Example 2.5, the ordering of rows indexed using the construction above are not necessarily the ordering as a subsequence of $(1,\,\ldots,\,p-1)$. Proposition 2.8 also implies that other than $\mathbbm{1}$, (p1)/2(p-1)/2 vectors in $\Phi_{p}$ have entries equal to $\zeta$ raised to some permutation of the set of the $\operatorname{QR}$ and (p1)/2(p-1)/2 vectors entries which have $\operatorname{NR}$ as the exponents.
We have one final lemma. As this will also be used to prove results about Paley ETFs $\Phi_{q}$ with $q$ a prime power, we state it in more generality than what is needed for this section.
@@
Let $q=p^{s}>3$ be a prime power with $q\equiv 3\bmod 4$. Let $X$ be a set of cardinality $q$. Assume that $G$ acts transitively and $2$-homogeneously but not $2$-transitively on $X$. Assume further that
$\tilde{G}=\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}$
is a subgroup of $G$, where
$\operatorname{A\Gamma L}(\mathbb{F}_{q})=\left\{\,f:\mathbb{F}_{q}\rightarrow\mathbb{F}_{q}\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\mathbb{F}_{q}^{\times},\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}.$
Then $G=\tilde{G}$.
Proof.
Theorem 1 of [Kan72] yields that any group which acts $2$-homogeneously but not $2$-transitively on a set of size at least $4$ must be a subgroup of $\operatorname{A\Gamma L}(\mathbb{F}_{q})$. Note that since $q\equiv 3\bmod 4$, $s$ and the cardinality of $\operatorname{QR}(\mathbb{F}_{q})$ must also both be odd, i.e., $\tilde{G}$ has odd order. Also, since $\operatorname{QR}(\mathbb{F}_{q})$ is index $2$ in $\mathbb{F}_{q}^{\times}$, $G$ is a maximal subgroup of $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ of odd order. It follows from the proof of Proposition 3.1 in [Kan69b] that a symmetry group which is $2$-homogeneous but not $2$-transitive must be of odd order. (Alternatively, one may note that $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ acts $2$-transitively on $\mathbb{F}_{q}$ and that $\tilde{G}$ is a maximal proper subgroup.) Hence, $G=\tilde{G}$, as desired. ∎
Note that when $q=p$, a prime, $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ is denoted $\operatorname{AGL}(\mathbb{F}_{p})$ and is parameterized by $\mathbb{F}_{p}$ and $\mathbb{F}_{p}^{\times}$ since there are no nontrivial field automorphisms.
Theorem 2.10.
$G:=\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}$
$\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{(p-1)/2}\cong\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,b\in\mathbb{F}_{p},\,\,m\in\operatorname{QR}(\mathbb{F}_{p})\,\right\}$
is the group of vector automorphisms of the Paley ETF $\Phi_{p}$. The vector and line symmetry groups are $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$. The symmetry groups act $2$-homogeneously but not $2$-transitively on the vectors, respectively the lines.
Proof.
We first show that $\{M^{k}T^{\ell}\}$ is isomorphic to the given subgroup of $\operatorname{AGL}(\mathbb{F}_{p})$ and that $\{M^{k}T^{\ell}\}\leq\operatorname*{Sym}_{v}(\Phi_{p})$. To that end, we note that the order of $M$ is $p$ and the order of $T$ is (p1)/2(p-1)/2. This also means that raising $M$ to elements of $\mathbb{F}_{p}$ is well-defined when interpreted mod $p$. Further, for $\alpha$ the primitive element used to define $M$, $k\in\{0,\,\ldots,\,p-1\}$, and $\ell\in\{0,\,\ldots,\,(p-3)/2\}$, we have by direct computation on the entries of the matrices
$T^{\ell}M^{k}T^{-\ell}=M^{k\alpha^{-2\ell}}.$ (1)
Clearly, $T^{\ell}\mathbbm{1}=\mathbbm{1}$ for any $\ell\in\{0,\,\ldots,\,(p-3)/2\}$. Thus, (1) yields for any frame vector $\varphi_{\alpha^{2k}}$ parameterized by a quadratic residue $\alpha^{2k}$,
$T^{\ell}\varphi_{\alpha^{2k}}=T^{\ell}M^{\alpha^{2k}}\mathbbm{1}=M^{\alpha^{2(k-\ell)}}T^{\ell}\mathbbm{1}=M^{\alpha^{2(k-\ell)}}\mathbbm{1}=\varphi_{\alpha^{2(k-\ell)}}.$
That is, $T^{\ell}$ maps vectors parameterized by quadratic residues to other vectors parameterized by quadratic residues cyclically. A similar calculation shows that left-multiplying by $T$ cyclically permutes the vectors parameterized by the quadratic non-residues. To be clear: $T$ acts on $\mathbb{C}^{(p-1)/2}$ by cyclically permuting the entries of vectors. When applied to $\Phi_{p}$, this cyclic permutation of the rows induces a cyclic permutation of the columns indexed by the quadratic residues and a cyclic permutation of the columns indexed by the quadratic non-residues. So, $\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}$ is a subgroup of the vector automorphisms. The commutation relation in (1) also yields
with corresponding group operation
[…]
which may be rewritten as
[…]
which is isomorphic to
$G:=\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,b\in\mathbb{F}_{p},\,\,m\in\operatorname{QR}(\mathbb{F}_{p})\,\right\}.$
Now we would like to show that $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$ and that the groups act $2$-homogeneously but not $2$-transitively on the vectors/lines.
Note that Proposition 2.8 implies that the vector symmetry group $\operatorname*{Sym}_{v}(\Phi_{p})$ and thus also the line symmetry group $\operatorname*{Sym}_{\ell}(\Phi_{p})$ act transitively since the vectors in $\Phi_{p}$ are the orbit of $\mathbbm{1}$ under $\{M^{k}\}$. We further have that $\{M^{k}T^{\ell}\}$ acts $2$-homogeneously on the vectors as follows. Let $x\neq y$ and $z\neq w$ be in $\mathbb{F}_{p}$. Since $p\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$. In particular either xyx-y or yxy-x is a quadratic residue, and the same is true for $z$ and $w$. Thus, without loss of generality, we pick an ordering so that $x-y,z-w\in\operatorname{QR}$. Then $M^{-y}\varphi_{y}=M^{-w}\varphi_{w}=\mathbbm{1}$, and $M^{-y}\varphi_{x}$ and $M^{-w}\varphi_{z}$ are both columns indexed by quadratic residues. We can then apply a translation $T^{\ell}$ as needed, which fixes $M^{-y}\varphi_{y}=M^{-w}\varphi_{w}=\mathbbm{1}$ and maps $M^{-w}\varphi_{z}$ to $M^{-y}\varphi_{x}$. Thus, $M^{y}T^{-\ell}M^{-w}$ maps $\{\varphi_{z},\varphi_{w}\}$ to $\{\varphi_{x},\varphi_{y}\}$.
We know from [IM24] that $\operatorname*{Sym}_{\ell}(\Phi_{p})$ is not $2$-transitive and thus neither is $\operatorname*{Sym}_{v}(\Phi_{p})$. It thus follows from Lemma 2.9 that $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$. ∎
This theorem (and the more general result Theorem 2.19) should not be incredibly surprising, given that Paley designs are the only $2$-homogeneous but not $2$-transitive Hadamard designs [Kan69b]. Note, however, that the Paley Hadamard designs built from $\mathbb{F}_{11}$ and $\mathbb{F}_{7}$ are $2$-transitive [Tod33, Kan69a], differentiating this result from the similar result for Paley ETFs. As we will see in Theorem 3.5, Paley skew-conference matrices, which are also built using quadratic residues, also yield an infinite class of ETFs which are $3$-homogeneous but not $3$-transitive.
2.3 Alternate Proof of Theorem 2.10
In this section, we will provide an alternate proof characterizing the symmetry groups of $\Phi_{p}$ which filters through the inner products, respectively, triple products (i.e., uses Theorem 1.14). This has the benefit that some of the techniques might be useful in the analysis of symmetry groups of other ETFs but the drawback that the explicit unitary matrices which are vector, respectively, line automorphisms are not constructed. We conclude with a short example of using the approach on another class of ETFs (Gabor-Steiner). In order to use Theorem 1.14, we need to characterize the inner products and triple products of the Paley ETF vectors.
Lemma 2.11.
@@
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Further let $\zeta$ be the primitive $p$th root of unity and $\alpha^{2}$ the generator of $\operatorname{QR}(\mathbb{F}_{p})$ used to produce $\Phi_{p}$. Finally, set
$a=\sum_{\ell=0}^{(p-3)/2}\zeta^{\alpha^{2\ell}}.$
@@
$\operatorname*{TP}(j,k,\ell)=\left\{\begin{array}[]{lr}a^{3};&j-k,k-\ell,\ell-j\in\operatorname{QR}\\[2.0pt] \left|a\right|^{2}a;&\textrm{ two of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \left|a\right|^{2}\overline{a};&\textrm{ one of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \overline{a}^{3};&\ j-k,k-\ell,\ell-j\in\operatorname{NR}\end{array}\right$.
@@
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\sum_{\ell=0}^{(p-3)/2}\zeta^{j\alpha^{2\ell}}\zeta^{-k\alpha^{2\ell}}=\sum_{\ell=0}^{(p-3)/2}\zeta^{(j-k)\alpha^{2\ell}}=\left\{\begin{array}[]{lr}(p-1)/2&j=k\\ a&j-k\in\operatorname{QR}\\ \overline{a}&j-k\in\operatorname{NR}\end{array}\right$.
since $(j-k)\alpha^{2\ell}\in\operatorname{QR}$ if and only if $j-k\in\operatorname{QR}$ and also $\operatorname{NR}=-\operatorname{QR}$. So, for distinct $j,k,\ell$, the desired result immediately follows. ∎
These triple products are distinct.
@@
Let $a$ be as in Lemma 2.11. Then $a^{3}$, $\left|a\right|^{2}a$, $\left|a\right|^{2}\overline{a}$, and $\overline{a}^{3}$ are pairwise distinct.